OpenStudy (anonymous):

solve 5x^2-y^2=0 3x^2+4y^2=0

OpenStudy (anonymous):

y^2=5x^2, then i get 100x^4+3x^2=0

OpenStudy (anonymous):

second isn't correct ! right hand shouldn't be zero.

OpenStudy (anonymous):

thats how its stated in the problem

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

sorry, I thought you typed it differently! \[y = \sqrt(5) x\]

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

and for the second part, \[y = 1/2 i \sqrt(3) x\]

OpenStudy (anonymous):

no, you doing it wrong

OpenStudy (anonymous):

what do you mean? that's the answer I'm pretty sure, or are you just trying to simplify?

OpenStudy (anonymous):

they are both circles with radius of 0 so isn't the answer (0,0)?

OpenStudy (anonymous):

yes that's the integer answer

OpenStudy (anonymous):

the answer for the variable y is what I posted

OpenStudy (anonymous):

no, you should solve for x and y, and you are solving it only for y

OpenStudy (anonymous):

then yes, x = 0 and y = 0

OpenStudy (anonymous):

alright